Midpoint/Distance/Linear Equations
Mario and Luigi need to get from their home to the race track in their hometown of coordinateville. They know that their home is located at (-4,-5) and the race track is at (-3,-7). Find the midpoint and the distance between their house and the race track. Also, what is the line of equation between their house and the race track?

Answer by nyc_function(2741)

Let M = midpoint

M = [(-3) + (-4)]/2,[(-5 + (-7)]/2

M = (-7/2, -12/2)

M = (-7/2, -6)....This is the midpoint.

=======================================

I guess by the line equation, you mean the equation of the line between the two given points.


In that case, we need to find the slope first.


Let little m = the slope


m = (-7 - (-5))/(-4 - (-3))


m = (-7 + 5)/(-4 + 3)


m = -2/-1

m = 2

The slope is 2.

We now use the point-slope formula and one of the given points to find your equation.

y - y1 = m(x - x1)....This is the point-slope formula.

y - (-5) = 2(x - (-4))

y + 5 = 2(x + 4)

y + 5 = 2x + 8

y = 2x + 8 - 5

y = 2x + 3